### v1

# FlowPolicy: Enabling Fast and Robust 3D Flow-based Policy via Consistency Flow Matching for Robot Manipulation

Abstract
Robots can acquire complex manipulation skills by learning policies from expert demonstrations, which is often known as vision-based imitation learning. Generating policies based on diffusion and flow matching models has been shown to be effective, particularly in robotics manipulation tasks. However, recursion-based approaches are often inference inefficient in working from noise distributions to policy distributions, posing a challenging trade-off between efficiency and quality. This motivates us to propose FlowPolicy, a novel framework for fast policy generation based on consistency flow matching and 3D vision. Our approach refines the flow dynamics by normalizing the self-consistency of the velocity field, enabling the model to derive task execution policies in a single inference step. Specifically, FlowPolicy conditions on the observed 3D point cloud, where consistency flow matching directly defines straight-line flows from different time states to the same action space, while simultaneously constraining their velocity values, that is, we approximate the trajectories from noise to robot actions by normalizing the self-consistency of the velocity field within the action space, thus improving the inference efficiency. We validate the effectiveness of FlowPolicy on Adroit and Metaworld, demonstrating a 7$\times$ increase in inference speed while maintaining competitive average success rates compared to state-of-the-art policy models. Codes will be made publicly available.
Introduction
Enabling robots to acquire diverse manipulation tasks by learning from expert demonstrations, commonly referred to as visual imitation learning, has been a long-standing objective in robotic learning and embodied AI (Atkeson and Schaal 1997; Argall et al. 2009; Wang et al. 2024a; Zare et al. 2024; Li et al. 2024). Prior work attempts to tackle this problem either by explicitly defining it as a supervised regression task (Zhang et al. 2018; Florence, Manuelli, and Tedrake 2019; Toyer et al. 2020; Rahmatizadeh et al. 2018), or by implicitly modeling action distributions using energy-based models (Florence et al. 2022; Jarrett, Bica, and van der Schaar 2020). The former learns the mapping from observations to actions, while the latter seeks to identify the minimal energy action through energy optimization. However, explicit regression approaches are often inadequate for capturing multi-modal demonstration behaviors and struggle with high-precision tasks (Florence et al. 2022). Conversely, energy-based models face challenges with training stability, primarily due to the necessity of negative sample extraction during the training process (Chi et al. 2023).
Figure 1: Comparison of FlowPolicy with the state-of-the-art method DP3 (Ze et al. 2024) and its lightweight version Simple DP3 in terms of inference time and average success rate on Adroit and Metaworld. Our FlowPolicy shows the best trade-off, achieving competitive results with an average inference time of 19.9ms. FlowPolicy is 7$\times$ faster than DP3 (19.9ms vs. 145.7ms) and 3$\times$ faster than Simple DP3 (19.9ms vs. 63.0ms).
Recent advancements in diffusion-based policy learning frameworks (Wang, Hunt, and Zhou 2022; Pearce et al. 2023; Reuss et al. 2023; Chi et al. 2023; Ze et al. 2024) have elevated this task to new heights, leveraging their powerful capability to model complex, high-dimensional robotic trajectories. Within this context, Diffusion Policy (Chi et al. 2023) defines policy learning as a conditional denoising diffusion process over the robot action space, conditioned on 2D observation features. The subsequent 3D Diffusion Policy (Ze et al. 2024), also known as DP3, further extends the conditional input to more complex 3D visual representations and is thus capable of acquiring generalizable visuomotor policies with only a few dozen expert demonstrations. Nevertheless, diffusion-based solutions are inevitably plagued by substantial runtime inefficiencies, as they typically require numerous sampling steps during inference to generate high-quality actions. While various sampling acceleration techniques (Song, Meng, and Ermon 2020; Song et al. 2023) have been proposed, the critical challenge of balancing efficiency and policy quality persists, severely limiting the practical application of these learned policies.
On the other hand, Flow Matching (FM) (Lipman et al. 2022), another class of generative model frameworks that directly defines probability paths via Ordinary Differential Equations (ODEs) to transform between noise and data samples, has demonstrated greater numerical stability and inference efficiency than diffusion families. The latest advancement in this area, Consistency Flow Matching (CFM) (Yang et al. 2024), further improves efficiency by directly defining a straight-line flow from any point in time to the same endpoint, enabling one-step generation for FM, which offers a potential direction for real-time policy generation. However, the conditional generation capability of CFM, particularly for embodied tasks involving complex 3D vision representations as conditions, remains underexplored.
In this paper, we address these challenges in policy generation by leveraging the concept of consistency flow matching, introducing a novel 3D flow-based framework for real-time robotic manipulation policy generation, named FlowPolicy. More specifically, we make the first attempt at conditional consistent flow matching in 3D robotic manipulation tasks. FlowPolicy allows for the fastest possible generation of target actions by directly defining straight-line flow conditions on the initial 3D point clouds. In addition, multi-segment training is applied to ensure the quality of action generation, aiming at a better trade-off between efficiency and effectiveness. We evaluate our framework on 37 robot manipulation tasks from Adroit and Metaworld, encompassing a wide range of tasks from simple to complex and dexterous manipulation. As illustrated in Fig. 1, our approach achieves a 7$\times$ reduction in average inference time while maintaining a competitive average success rate compared to state-of-the-art methods based on 3D visual conditional representation. By combining consistency flow matching with 3D condition-based robot manipulation, We enable the real-time decoding of high-quality robot actions with on-step inference.
Figure 2: Overall pipeline. The above section is a visualization of FlowPolicy. By defining a straight-line flow the data can flow the fastest from the noise distribution to the action distribution (Adroit: Open the door). The following section shows the details of FlowPolicy. Given a certain number of expert presentations, it is first converted into 3D point clouds. The 3D point clouds and the robot state are then fed into two encoders to obtain the compact 3D visual representation and the robot state embedding, respectively. Finally, a straight-line flow is learned by conditional consistency flow matching to generate high-quality robot actions and perform the corresponding tasks (Metaworld: Assembly) at real-time inference speed.
In summary, our main contributions are threefold:
  • We first propose a 3D flow-based policy generation framework that conditions the 3D visual representation and can generate robust robotic actions with few demonstrations, namely FlowPolicy.
  • The proposed FlowPolicy significantly narrows the gap between efficiency and quality in vision-based imitation learning algorithms, enabling real-time policy generation across diverse manipulation tasks.
  • We conduct extensive experiments on 37 simulation tasks from Adroit and Metaworld, demonstrating that our approach achieves state-of-the-art average success rates.
Related Work
Application of Diffusion Models to 3D Robotics
Diffusion models are a family of generative models that are highly generalizable and easy to train, which are initially utilized for image generation tasks (Batzolis et al. 2021; Epstein et al. 2023). As a subversive approach, diffusion models progressively refine Gaussian noise by recursively refining it to the point where it satisfies the data distribution. Recently, it has been introduced to the field of robotics, including reinforcement learning (Römer et al. 2024), imitation learning (Wang et al. 2024b), trajectory Optimization (Carvalho et al. 2023) and grasping (Urain et al. 2023), and has shown impressive performance. Denoising Diffusion Probabilistic Models (DDPM) (Ho, Jain, and Abbeel 2020) generates a clean sample from noise by solving a Stochastic Differential Equation (SDE). However, numerical simulations of diffusion processes typically involve hundreds of steps, leading to significant slowness. Denoising Diffusion Implicit Modeling (DDIM) (Song, Meng, and Ermon 2020) speeds up inference by solving the inverse process as an Ordinary Differential Equation (ODE), but this sacrifices generation quality. Although adaptive ODE solvers exist, they do not significantly reduce the number of inference steps. In 3D-related robot manipulation tasks with high-dimensional control, it is challenging to achieve real-time inference speed with high-quality actions based on diffusion models. The main obstacles arise from the uncertainty of the diffusion process and its recursive nature.
To improve the efficiency problem of the diffusion process, Song (Song et al. 2023) proposes the consistency model. The consistency model is a novel model that directly maps noise to a data distribution by learning to map points at arbitrary moments on the probabilistic flow ODE (PF-ODE) trajectory to the origin (i.e., clean data) to support one-step or several-step generation. It has been utilized in several domains such as high fidelity image generation (Luo et al. 2023), text (Seo et al. 2023) and video (Chai et al. 2023) due to its efficient expressiveness. While consistency models support one-step generation, distillation is usually required to reach the desired results. This results in its limited application in robotic manipulations.
Flow Matching
Flow Matching (FM) (Lipman et al. 2022) is an emerging family of generative models based on optimal transport theory (Mozaffari et al. 2016), which aims to estimate a vector field that yields an ODE and admit the desired probabilistic path. FM allows direct regression to the goal of estimating the vector field, and hence it is simpler and more numerically stable than DDPM and DDIM. By avoiding estimating noise and instead matching a path from the noise to the target, FM enables faster inference, which is crucial in real-time robot manipulations. Previous works have demonstrated the effectiveness of FM for inverse reinforcement learning (Chang et al. 2022), multi-support motion (Rouxel et al. 2024), and robot deployment (Feng et al. 2023). However, FM-based generation of complex robot action has rarely been reported, especially involving complex 3D spatial representations. In this work, we explore the feasibility of applying FM to 3D robot manipulation and hope to achieve real-time generation through consistency constraints.
Recently, Yang (Yang et al. 2024) proposed to perform consistency in velocity field space rather than sample space to improve the flow matching process (CFM). CFM directly defines a straight-line flow from any time to the same endpoint and imposes constraints on its velocity values. Also, the multi-segment training technique allows CFM to achieve a better trade-off between sampling quality and velocity. Therefore, CFM is an elegant balance of efficiency and performance. More importantly, Yang  (Yang et al. 2024) demonstrated that CFM can be trained to produce a robust flow model without the aid of distillation, which is valuable to robots performing unseen tasks, as it is difficult to find a robust teacher model in practical applications. While consistency-based models have shown promising results in some areas of embodied intelligence (Ding and Jin 2024; Chen, Li, and Zhao 2024; Prasad et al. 2024), its effectiveness has not been demonstrated on embodied tasks with complex 3D vision representation as conditions, especially for flow-based methods. To address this issue, we propose FlowPolicy, a real-time 3D policy generation framework based on consistency flow matching.
Method
We focus on imitation learning to enable robots to acquire complex manipulation skills. Specifically, we expect a limited number of expert demonstrations to teach an agent to learn a policy $\pi:O\Longrightarrow A$, i.e., mapping actions from visual observations $\textit{o}\in{O}$ to actions $\textit{a}\in{A}$. Visual observations include the robot state and scene point clouds, and actions are usually sequences of trajectories of the robot to accomplish a specific task. Diffusion-based and flow-based methods can generate high-quality robot actions, but expensive iterative sampling makes them unable to accomplish real-time operations, and thus the application scenarios are very limited. While CFM balances efficiency and effectiveness, few existing works have explored the application of 3D condition generation based on consistent flow matching to robotic manipulation. Therefore, we propose FlowPolicy, a conditional consistency flow matching model, which guarantees the generation of high-quality actions while also accomplishing one-step inference for real-time applications. This section is organized as follows, we first review flow matching and consistent flow matching, and then introduce the pipeline of FlowPolicy. Finally, we describe the design details of each component.
Consistency Flow Matching
We first introduce the notion of flow matching (Lipman et al. 2022), given a ground truth vector field $u(t,x)$ that generates the probabilistic path $p_{t}$ under two marginal constraints, $p_{t=0}=p_{0}$ and $p_{t=1}=p_{1}$. The expectation is to find an ODE whose solution transmits the noise ${x_{0}}\sim{p_{0}}$ to the data ${x_{1}}\sim{p_{1}}$:
$\left\{\begin{array}[]{rcl}\frac{d{\xi}_{x}(t)}{dt}={\nu}_{\theta}(t,{\xi}_{x}(t))\\ {\xi}_{x}(0)=x\end{array}\right.$ (1)
${\xi}_{x}(t)$ is called a flow. The goal of flow matching is therefore to learn a vector field ${\nu}_{\theta}(t,x_{t})$:
$L_{\text{FM}}(\theta)=E_{t,p_{t}}{\|{{\nu}_{\theta}(t,x_{t})-u(t,x_{t})}\|}^{2}_{2}$ (2)
Flow matching is a simple and attractive objective. However, lack of a prior knowledge about $u$ and $p_{t}$, conditional flow matching(Lipman et al. 2022) is typically employed to regress ${\nu}_{\theta}(t,x_{t})$ given a conditional vector field $u(t,x_{t}|{x_{1}})$ and a conditional probability path $p(t,x_{t}|{x_{1}})$:
$L_{\text{FM}^{\ast}}(\theta)=E_{t,q_{(}{x_{1}})}E_{t,p_{t}|{x_{1}}}{\|{{\nu}_{\theta}(t,x_{t})-u(t,x_{t}|x_{1})}\|}^{2}_{2}$ (3)
Learning straight-line flows enables faster inference efficiency. Consistency flow matching (CFM)  (Yang et al. 2024) is a generalized method for efficiently learning straight-line flows without approximating the entire probabilistic path. Consistent velocity is defined as ${\nu}(t,{\xi}_{x}(t))={\nu}(0,x)$, indicating that the velocity along the solution of Equation 1 remains constant. Due to the complexity of the target distribution solution, CFM does not regress directly on the ground truth vector field, instead, it directly defines a velocity-consistent straight-line flow that departs at different times to the same endpoint. CFM can regress on the following objective to learn a velocity-consistent vector field:
${L_{\theta}=E_{t{\sim}\mu}E_{x_{t},x_{t}+\Delta{t}}\|f_{\theta}(t,x_{t})-f_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}_{2}}\\ {+\alpha{\|{\nu}_{\theta}(t,x_{t})-{\nu}_{\theta^{-}}(t+\Delta{t})\|^{2}_{2}}}\\ f_{\theta}(t,x_{t})=x_{t}+(1-t)*{\nu}_{\theta}(t,x_{t})$ (4)
where $\mu$ is a uniform distribution on $[0,1-\Delta{t}]$, $\alpha$ is a positive scalar, and $\Delta{t}$ is the time interval, a small and positive scalar. ${\theta}^{-}$ denotes the running average of past $\theta$ values computed using an exponential moving average (EMA). Training $L_{\theta}$ not only regularizes the velocity but also learns the data distribution.
Additionally, CFM allows for multi-segment training to enhance its ability to express the distribution of transmission through $p_{t}$. Specifically, multi-segment CFM splits time into uniform subsets, normalizes a consistent vector field ${\nu}(t,x_{t})$ within each subset, and in turn constructs segmented linear trajectories. This approach relaxes the velocity consistency requirement and makes data transmission more flexible and efficient. Given the number of segments K, the training loss of segmented CFM is as follows:
$L_{\theta}=E_{t{\sim}{\mu}^{i}}{\lambda}^{i}E_{x_{t},x_{t}+\Delta{t}}\|f^{i}_{\theta}(t,x_{t})\\ -f^{i}_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}{{}_{2}}\\ {+\alpha{\|{\nu}^{i}_{\theta}(t,x_{t})-{\nu}^{i}_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}{{}_{2}}}}\\ f^{i}_{\theta}(t,x_{t})=x_{t}+(i+1)/k-t*{\nu}^{i}_{\theta}(t,x_{t})$ (5)
Here, ${\mu}^{i}$ is a uniform distribution on $[i/K,(i+1)/K-\Delta{t}]$. The moving step of each segment is controlled by a very small positive number $\Delta{t}$. $x$ and $x_{t}$ both follow a well-established distribution pattern. The flow and vector fields of each segment are described by ${\nu}^{i}_{\theta}(t,x_{t})$, which depend on the parameters $\theta$ and time $t$. In addition, ${\lambda}^{i}$ is a positive weighting factor used to balance the importance of the different sections, especially since vector fields in the middle section are usually more difficult to train.
FlowPolicy
The overall pipeline of our proposed FlowPolicy is presented in Fig. 2. FlowPolicy takes the powerful 3D diffusion policy (DP3) as the underlying architecture. DP3 is the state-of-the-art 3D vision-based conditional diffusion policy model. We address the inefficiency of 3D policy inference by constructing conditional consistency flow matching as a policy generator while ensuring the quality of task completion. Specifically, we consider the observed 3D point cloud as a visual conditional representation, which is combined with state embedding to guide the flow model to learn the manipulation policy. In the training phase, we regress the consistency vector field and learn the straight-line flow through velocity consistency loss, which directly generates the robot action sequences. Depending on the consistency training and condition representation, FlowPolicy can generate high-quality actions in one-step inference and accomplish complex manipulation tasks.
Methods NFE Adroit Metaworld Average
Hammer Door Pen Easy(21) Medium(4) Hard(4) Very Hard(5)
DP3 10 146.6$\pm$0.7 146.8$\pm$2.0 144.8$\pm$2.9 145.9$\pm$2.1 144.7$\pm$2.3 148.6$\pm$3.4 143.2$\pm$2.6 145.7$\pm$2.3
Simple DP3 10 64.5$\pm$1.4 64.2$\pm$2.4 71.2$\pm$9.3 62.7$\pm$6.2 59.4$\pm$4.1 63.0$\pm$6.8 65.3$\pm$6.7 63.0$\pm$5.9
FlowPolicy (Ours) 1 20.4$\pm$0.5 19.3$\pm$0.9 20.6$\pm$0.2 19.9$\pm$0.1 19.8$\pm$0.3 20.1$\pm$0.2 20.0$\pm$0.1 19.9$\pm$0.2
Table 1: Quantitative comparison of runtime between state-of-the-art policy models and the proposed FlowPolicy. We evaluate 3 dexterity tasks from Adroit and 34 challenging tasks from Metaworld across 3 random seeds and report inference time per step (ms) with standard deviation. Our FlowPolicy enables real-time robot operations with an average time of 19.9ms in one-step inference, which is 7 $\times$ faster than DP3 and 3 $\times$ faster than its lightweight version, SimpleDP3.
Methods NFE Adroit Metaworld Average
Hammer Door Pen Easy(21) Medium(4) Hard(4) Very Hard(5)
DP3 10 100$\pm$0 56$\pm$5 46$\pm$10 87.3$\pm$2.2 44.5$\pm$8.7 32.7$\pm$7.7 39.4$\pm$9.0 68.7$\pm$4.7
Simple DP3 10 98$\pm$2 40$\pm$17 36$\pm$4 86.8$\pm$2.3 42.0$\pm$6.5 38.7$\pm$7.5 35.0$\pm$11.6 67.4$\pm$5.0
FlowPolicy (Ours) 1 100$\pm$0 58$\pm$5 53$\pm$12 90.2$\pm$2.8 47.5$\pm$7.7 37.2$\pm$7.2 36.6$\pm$7.6 70.0$\pm$4.7
Table 2: Comparisons on success rate between state-of-the-art policy models and the proposed FlowPolicy. We evaluate 3 dexterity tasks from Adroit and 34 challenging tasks from Metaworld across 3 random seeds and report the success rate (%) with standard deviation. For fair comparisons, DP3 and Simple DP3 are reproduced under the same experimental environment and expert demonstrations as ours. Our method achieves competitive success rates compared to the current state-of-the-art approach DP3 while enabling real-time, one-step inference.
3D visual conditional representation
FlowPolicy applies 3D point clouds as a guide to generate execution-specific policies given its rich spatial information. Specifically, we first acquire single-view depth-observation images and then convert the depth information into a point cloud via camera intrinsics and extrinsic. The single-view image is more practically applicable and the point cloud provides a richer spatial representation. Further, a farthest point sampling (FPS) (Qi et al. 2017) is applied to downsample the point cloud to preserve spatial wholeness and reduce redundancy. Finally, we employ a lightweight MLP to encode the point cloud into a compact 3D representation as the visual conditional representation for flow-based policy. With the conditional constraints of the point cloud data, FlowPolicy enables the efficient capture of spatial information that is critical to fine robotic manipulation tasks.
Conditional Trajectory Generation with FlowPolicy
For FlowPolicy, given the following definitions:
$a^{\text{src}}{\sim}\Gamma^{\text{src}},a^{\text{tar}}{\sim}\Gamma^{\text{tar}}({\cdot}|s,v),t{\sim}\mu[\epsilon,1-\epsilon]$ (6)
where $\Gamma^{\text{src}}$ is the source distribution, usually described as Gaussian noise $\Gamma^{\text{src}}=N(0,I)$. $\Gamma^{\text{tar}}$ is the target distribution conditioned by the state $s$ and visual representation $v$. $a^{\text{src}}$ and $a^{\text{tar}}$ are denoted as sequences of trajectories sampled from the source and target distributions. $t$ obeys a uniform distribution $t\in[\epsilon,1-\epsilon]$. Define $a_{t}$ to be the interpolated trajectory between the source and destination trajectories at the transmission time step $t$. $F$ is the flow model conditioned by the state $s$ and visual representation $v$, FlowPolicy can be simply described as:
$\text{FlowPolicy}\ F:a,t,s,v\longrightarrow{\Delta{a}}$ (7)
To train a high-quality one-step generator, we employ a two-segment CFM and use backpropagation to minimize the following equation:
$L_{\text{FlowPolicy}}(\theta)=E_{t{\sim}{\mu}^{i}}{\lambda}^{i}E_{a_{t},a_{t}+\Delta{t}}\|f^{i}_{\theta}(t,a_{t},s,v)\\ -f^{i}_{\theta^{-}}(t+\Delta{t},a_{t+\Delta{t}},s,v)\|^{2}_{2}\\ {+\alpha{\|{\nu}^{i}_{\theta}(t,a_{t},s,v)-{\nu}^{i}_{\theta^{-}}(t+\Delta{t},a_{t+\Delta{t}},s,v)\|^{2}_{2}}}\\ f^{i}_{\theta}(t,a_{t},s,v)=a_{t}+(i+1)/k-t*{\nu}^{i}_{\theta}(t,a_{t},s,v)$ (8)
By training the two-segment loss above two consistent vector fields can be learned, which in turn constructs a two-segment linear trajectory that transforms the noisy trajectory into the target trajectory. In the inference process, the noise trajectories are sampled from the source distribution $a_{0}\in a^{\text{src}}$ and then predicted by a flow model to obtain action trajectories $a_{1}\in a^{\text{tar}}$.
FlowPolicy provides a framework for generating high-quality actions in a one-step process that promotes a trade-off between efficiency and effectiveness. Compared to DP3 and Simple DP3, FlowPolicy utilizes minimal inference time to achieve a competitive average success rate.
Figure 3: Qualitative Comparison of FlowPolicy and DP3 (Ze et al. 2024) on two challenging manipulation tasks from Adroit and Metaworld. Our method successfully generates high-quality actions at real-time speeds, completing these tasks effectively, whereas DP3 either produces lower-quality actions (left) or fails to complete the task (right).
Experiments
Dataset and Implementation Details
Simulation Benchmarks
We have chosen two preeminent environmental simulators, Adroit (Rajeswaran et al. 2018) and Metaworld (Yu et al. 2021), as our benchmarks for a comprehensive set of 37 tasks. The Adroit benchmark presents three challenges of dexterous manipulation tasks, where the execution of complex tasks is facilitated by the intricate control of a high-dimensional, articulated hand. In contrast, the Metaworld benchmark offers a diverse array of tasks that span the spectrum of difficulty levels, from easy to very hard, typically surmounted with the aid of a two-fingered robotic arm. For the procurement of expert demonstrations, we have harnessed a well-trained heuristic strategy, designed to yield datasets of exceptional quality.
Baselines
The main objective of this work is to leverage 3D point clouds as visual conditional representations, aiming to elevate the robot manipulation to the real-time level via a consistency flow matching strategy. Therefore, we focus on comparing DP3, a state-of-the-art 3D conditional diffusion-based strategy model. In addition, its lightweight version, simple DP3, is also compared. For a fair comparison, the DP3 and Simple DP3 parameter settings in our experimental setup are the same as in (Ze et al. 2024).
Evaluation Metrics
We evaluate 20 episodes per 200 training epochs and record the mean of the five highest success rates. To avoid chance error, each task is run repeatedly under three different random seeds and their means and variances as well as inference times are calculated. In addition, the Number of Function Evaluations (NFE) (Yang et al. 2024) is evaluated as a metric of inference efficiency. The higher the success rate, the better, and the lower the NFE and inference time, the better.
Implementation Details
For each task, we generate a compact conditional embedding consisting of a 3D point cloud and the robot’s state with a dimension of 64. Additionally, our model employs an observation horizon of two steps, signifying that it leverages the point clouds from the two most recent time frames as conditional inputs for policy generation. Focused on efficient imitation learning, we assign ten expert demonstrations to each task. To mitigate unnecessary computational costs, the image size is cropped to a resolution of 84$\times$84 pixels and the point cloud is downsampled to 512 points for Adroit or 1024 points for Metaworld using the farthest point sampling method. During the training phase, the weights are updated using the AdamW optimizer with a learning rate of 1e-4, and a batch size of 128. The training process is iterated for 3,000 epochs. EMA decay rate is set to the default value of 0.95. Additionally, both action and state inputs are normalized to the interval [-1, 1] to facilitate the training process, with actions being unnormalized before execution. Our model and all compared methods are developed based on the PyTorch framework and evaluated on a single NVIDIA RTX 2080Ti GPU.
Comparison with State-of-the-art Methods
Quantitative Comparisons on Runtime
Table 1 reports the average inference time of each model for 3 dexterity tasks from Adroit and 34 different difficulty tasks from Metaworld. Our FlowPolicy shows impressive results by reducing the average inference time to 19.9ms, an improvement of more than 7$\times$ (19.9ms vs. 145.7ms) compared to the DP3 model. Even for the lightweight version of Simple DP3, FlowPolicy improves the inference efficiency by more than 3$\times$ (19.9ms vs. 63.0ms). Since a consistent velocity field provides a more direct linear path result in generating actions faster and more efficiently. More interestingly, our approach has less variation in inference time. Specifically, FlowPolicy has less fluctuation in the average runtime of performing individual tasks and overall tasks, with a potential explanation attributed to the deterministic sampling policy of consistent flow matching. The results show that faster ODE trajectories can be determined and the inference efficiency can be improved by directly defining a straight-line flow, which is crucial for realizing real-time robot operation.
Figure 4: Illustrations of the learning curves. Compared to Simple DP3 and DP3, FlowPolicy demonstrates higher stability, learning efficiency, and success rates.
Quantitative Comparisons on Success Rate
[…]
Figure 5: Ablation on the number of expert demonstrations. We choose four typical tasks to explore the impact of different numbers of demonstrations on FlowPolicy and DP3. Both generally improve the accuracy with more demonstrations, but FlowPolicy typically has a higher success rate and avoids the performance bottleneck as presented in DP3.
Qualitative Comparisons on Manipulation Tasks
As illustrated in Figure 3, we selected two typical qualitative examples in Adroit and Metaworld, generated by DP3 and FlowPolicy, respectively. From left to right, this sequence delineates the holistic progression of the task’s execution, with each frame of the sequence corresponding to a distinct phase in the operation, as performed by the agent. On the left is a pen-turning task involving complex dexterity manipulation, which requires turning the blue-capped pen in the dexterous hand to a specific position (i.e., the green-capped pen serves as the reference angle.). Although DP3 accomplishes the dexterity task, the diffusion policy generated based on DP3 fails to ensure consistency with the target pen in 3D space compared to FlowPolicy. Specifically, DP3 turns one end of the pen to a state so high that it is darker in the camera view. In contrast, FlowPolicy generates higher-quality action sequences to turn the pen to the most appropriate angle. More importantly, FlowPolicy gets closer to the target position faster at the same frame. The right case is a difficult ‘Pick-Place’ task in Metaworld, which requires the use of parallel gripping clips to carry a red cube to a blue target location. DP3 unsuccessfully picks up the red cube and fails the task. Instead, FlowPolicy correctly locates the position of the red cube and successfully carries it to the target location. The above findings show that our FlowPolicy not only accomplishes the tasks that DP3 cannot accomplish well, but also achieves a higher quality and speed of accomplishment than DP3. This further demonstrates that our proposed FlowPolicy has a more delicate understanding and generative ability in robot manipulation tasks.
Analysis of Learning Curves
Similarly, we selected four representative tasks in Adroit and Metaworld and visualized their learning curves, as illustrated in Figure 4. Although we defined 3000 epochs for each task, most of the tasks converged and leveled off at 500–1000 epochs. For the simple-level task ‘button press wall’, all three models learned the correct policy easily and converged with a small number of iterations. Although Simple DP3 and DP3 have high success rates, their success rates fluctuate as training progresses. On the contrary, our FlowPolicy learns the button policy well and maintains it, demonstrating the stability of FlowPolicy. For another simple level task ’reach wall’ and a medium difficulty task ‘hammer’, all three models learn faster, but Simple DP3 and DP3 reach a performance bottleneck, while our method achieves a a higher task success rate. The learning curves for the dexterity task in Adroit show similar conclusions. The above results show that FlowPolicy can learn policies faster and better, and demonstrates certain stability.
Ablation Studies on Expert Demonstrations
The success rate of the agent in accomplishing tasks depends on the number and quality of expert demonstrations, where the quality of demonstrations is ensured by a heuristic algorithm. We further conduct ablation studies to verify the influence of the number of expert demonstrations. As illustrated in Figure 5, the success rate variation curves for typical tasks with different number of demonstrations. For non-hard tasks (i.e., ‘Reach Wall’, ‘Hammer’), FlowPolicy can outperform DP3 with a limited number of presentations. For hard-level tasks (i.e., ‘Pick-Place’), the success rate of the task can be significantly improved by increasing the number of expert presentations, as shown in both FlowPolicy and DP3. However, DP3 reaches performance saturation more easily, which even leads to a decrease in the success rate when the number of demonstrations is sufficient (i.e., ‘Coffee Pull’), while FlowPolicy performs better. The experiments reveal a better trade-off between the number of expert demonstrations and the success rate of the proposed FlowPolicy, i.e., the ability to learn effective data distributions from a small amount of data and to break through performance bottlenecks when the amount of data is sufficient.
Conclusion
In this work, we have introduced FlowPolicy, a novel 3D flow-based visual imitation learning algorithm. It can be summarized as a policy-generating model based on 3D point cloud and consistency flow matching. For the first time, we have explored the challenge of conditional generation within consistency flow matching, particularly under complex 3D visual representations. FlowPolicy accelerates the transfer of data from the noise to the action space by defining straight-line flows, thereby improving the inference efficiency of robotic manipulation tasks at the real-time level. Evaluations across 37 tasks of diverse challenges on Adroit and Metaworld platforms demonstrate that FlowPolicy achieves competitive success rates with real-time inference speed. We hope our work will advance 3D visual-based imitation learning towards practical applications.